OpenStudy (anonymous):

One number is five more than another, and their sum is three less than three times the smaller. Find the numbers. If x represents the smaller number, which equation could be used to solve for x?

OpenStudy (anonymous):

one number is x assume that the other number is y since x is the smaller number, and one number is 5 more than the other therefore, x=y+5 (equation 1) and their sum (x+y)is 3 less than 3 times the smaller(3x-3) i.e. x+y=3x-3 which gives 2x-y=3 (equation 2) you can solve the to equations to find x

OpenStudy (freckles):

6 is 5 more than 1 (6=5+1) but 6>1 so I think the first equation should be y=x+5 since you said x is the smaller of x and y

OpenStudy (anonymous):

oh yeah sry ... y=x+5